archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    CLOSED Percent Complete Details We have a VDSL card on DSLAM RNY93 with faulty ports Date Wednesday 08 October 2014 15 13PM Reason for closing Done Comments 2 Comment by OVH Wednesday 14 May 2014 08 37AM We reset the

    Original URL path: http://status.ovh.es/?do=details&id=6893 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    Backend Core Status CLOSED Percent Complete Details We have a new Paris gateway for the national France Télécom collect It will provide redundancy in relation to the Roubaix gatewayn for the national collect We will also change the configuration of devices to evenly distribute the national collect customers on the two gateways of Roubaix and Paris The PPP sessions will be reset on April 30th at 4 30 to reflect

    Original URL path: http://status.ovh.es/?do=details&id=6822 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Status CLOSED Percent Complete Details An optical intervention will take place between Lormont LOR33 and Ambares AMB33 The Ambares site is isolated Date Wednesday 23 April 2014 12 39PM Reason for closing Done Comments 1 Comment by OVH Wednesday 23

    Original URL path: http://status.ovh.es/?do=details&id=6784 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    and the optic device on the SL869 site in the Lyon area are no longer responding We will investigate Date Wednesday 23 April 2014 04 20AM Reason for closing Done Comments 2 Comment by OVH Tuesday 22 April 2014 17 20PM There was a electrical power cut on the SL869 site up 2 min 1 user load average 0 34 0 20 0 08 The connections should return The DSLAM

    Original URL path: http://status.ovh.es/?do=details&id=6783 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Maintenance Category Backend Core Status CLOSED Percent Complete Details We currently have issues in displaying ADSL access on manager We will investigate Date Wednesday 08 October 2014 15 13PM Reason for closing Done Comments 1 Comment by OVH Sunday 20

    Original URL path: http://status.ovh.es/?do=details&id=6771 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Complete Details We detected a fiber cut off on the loop Paris Couronne a ticket was opened The fiber is folded an intervention is in progress on site NRA CAC94 will be inaccessible during the intervention We expect to resume

    Original URL path: http://status.ovh.es/?do=details&id=6704 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Status CLOSED Percent Complete Details There has been a loss of traffic on the Roubaix LNS with the FT collect The PPP sessions were reset Date Tuesday 08 April 2014 17 26PM Reason for closing Done Comments 1 Comment by

    Original URL path: http://status.ovh.es/?do=details&id=6688 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Incident Category Backend Core Status CLOSED Percent Complete Details We have detected a cut in the fibre on the Paris petite couronne loop a ticket has been opened Securisation will take over Date Tuesday 08 April 2014 04 58AM Reason

    Original URL path: http://status.ovh.es/?do=details&id=6676 (2015-10-15)
    Open archived version from archive



  •